Hi I'm Angel
Please Answer
The Sum of two numbers is 44. If the larger number is two less than three times the smaller number, find the number.

please answer thank you!

Answer by ewatrrr(24785)

Hi,
Let x represent the smaller number, then (3x-2) is the larger number
.
Question states
x + (3x-2)= 44
.
simplify and slove
4x - 2 = 44
4x = 46
x = 11 1/2 is the smaller number (32 1/2 the larger)
.
check your answer
11.5 + 32.5 = 44